OpenStudy (anonymous):

The point (5, -1) is a solution of the equation: y = 2x − 11 true or false?

OpenStudy (igreen):

Plug in 5 for x and -1 for y. y = 2x - 11 -1 = 2(5) - 11 -1 = 10 - 11 Can you subtract 10 - 11?

OpenStudy (anonymous):

yeah..?

OpenStudy (igreen):

What is it?

OpenStudy (anonymous):

-1, so its.. false?

OpenStudy (igreen):

No, that gives us -1 = -1, so (5, -1) is a solution.

OpenStudy (anonymous):

oh okay i get it now. thanks!

OpenStudy (igreen):

No problem.